# Interacting with Thoughtful AI

### v1

@@ Abstract.
We envision the concept of Thoughtful AI, a new human–AI interaction paradigm in which the AI behaves as a continuously thinking entity. Unlike conventional AI systems that operate on a turn-based, input-output model, Thoughtful AI autonomously generates, iterates, and communicates its evolving thought process throughout an interaction. In this position paper, we argue that this thoughtfulness unlocks new possibilities for human-AI interaction by enabling proactive AI behavior, facilitating continuous cognitive alignment with users, and fostering more dynamic interaction experiences. We outline the conceptual foundations of Thoughtful AI, illustrate its potential through example projects, and envision how this paradigm can transform human–AI interaction in the future.
@@ 1. Introduction
In 1950, Alan Turing famously asked, “Can machines think?” (Turing, 1950) This question has has inspired decades of research in artificial intelligence (AI), with efforts ranging from symbolic reasoning and expert systems to today’s large language models (LLMs). Early philosophical and computational models of cognition viewed thought as a defining characteristic of intelligence—whether in humans or machines (Descartes, 1996; Newell, 1972; Miller and Cohen, 2001; Raichle et al., 2001). Recent LLMs have reignited discussions about what it means for an AI to “think”. Techniques such as Chain-of-Thought prompting (Wei et al., 2022) demonstrate that generating intermediate reasoning steps can significantly enhance model performance on complex tasks. Building on this foundation, models like OpenAI o1 (Jaech et al., 2024), Gemini Flash Thinking (Google DeepMind, 2025), and DeepSeek R1 (Guo et al., 2025) integrate reasoning into their training processes. These developments mark substantial progress in AI’s ability to problem-solve.
However, despite these advancements, current AI interaction paradigms remain fundamentally constrained. Existing systems still operate within a turn-based, input-output framework: users issue a prompt, the AI generates a single response, and the cycle repeats. This limitation becomes evident in everyday AI interactions. Imagine a user asking ChatGPT: “Help me plan a surprise birthday party for my friend.” ChatGPT would generate a static response—perhaps listing general steps like choosing a venue, sending invitations, and selecting a menu. At this point, the AI has already stopped processing and just passively waits for the user’s next input. If the user later realizes they need a budget-friendly vegan restaurant for six people in New York, they must start a new turn, manually refining their query.
In contrast, consider instead how humans naturally collaborate during a brainstorming session in front of a whiteboard. Each participant holds an internal train of thought that continuously updates in the background, even while someone else is speaking. In parallel, they may externalize immature, intermediate thoughts by sketching or annotating on the whiteboard, inviting immediate refinements and comments from others. This thinking process does not pause in discrete cycles: conversation flows as people update, erase, and reorganize ideas in real time. Meanwhile, private reflections that are not yet shared continue to shape what eventually gets added to the board.
@@
At the core of Thoughtful AI is the notion that AI’s thought is not merely hidden backend computation steps but a fundamental modality that contribute to AI’s capabilities to interact with humans. We identify four key traits that define this paradigm:
@@ 2. What makes AI “Thoughtful”?
Discussions around “thought” in LLM research often focus on a chain-of-thought approach, where a model generates intermediate reasoning steps to improve performance on reasoning tasks. From an HCI perspective, however, thought can be understood more broadly as a system’s internal process that evolves during an interaction and selectively surfaces information to users in real time. We define Thoughtfulness as:
@@
Thoughtfulness of AI refers to a system’s ability to continuously generate, develop, and selectively communicate its intermediate processes and responses over the course of an interaction.
@@ 2.1.1. Intermediate Medium
In most existing AI systems, users see only a final answer or recommendation. This one-shot approach is limiting: once the user receives an answer, any clarifications or follow-up requests often require multiple back-and-forth queries. If the AI’s thinking was flawed or incomplete, the user has no window into how or why the system arrived at its conclusion.
@@ 2.2. Implications for HCI
The four traits of Thoughtful AI—an intermediate medium, full-duplex process, intrinsic driver, and a shared cognitive space—open up new possibilities for how humans interact with AI. Below, we discuss three implications for Human-AI Interaction.
@@ 2.2.2. Continuous Cognitive Alignment
2.2.3. Beyond Chat and Turn-based Interaction
Finally, Thoughtful AI prompts us to reconsider the fundamental structure of human–AI dialogue. Traditional chatbot interfaces operate in discrete, back-and-forth messages, with the AI effectively going idle after sending each response. By contrast, a full-duplex process (Trait 2) enables the AI to continuously think and listen, even when the user is not actively providing input. This opens the door to interaction models that transcend simple chat boxes. For instance, an AI planning assistant could silently update its suggestions as it overhears new constraints in a virtual meeting, intervening only when necessary (Traits 2 and 3). More interestingly, AI and users can build a shared cognitive space (Trait 4), adding thoughts, annotations, and partial ideas to a collective interface. Rather than a linear log of exchanges, this collaborative canvas allows ideas to branch, merge, and evolve continuously—resembling a dynamic mind-map or sketchnote more than a turn-based chat.
@@ 3. Example Projects
To further illustrate the practical applications of Thoughtful AI, we present two projects that embody its core principles. The first, Inner Thoughts, explores how AI can proactively engage in conversations by developing and evaluating its internal thoughts before participating. The second, Interactive Thoughts, investigates AI-generated thoughts as an interface component that allows users to observe and interact with in real-time.
@@ 3.1. Inner Thoughts
3.2. Interactive Thoughts
While Inner Thoughts explores AI’s ability to think internally, Interactive Thoughts investigates a complementary question: what if users could actively observe and engage with AI’s thought process?
@@ 3.2. Interactive Thoughts
By contrast, Interactive Thoughts envisions a system where users can engage with AI-generated ideas before they fully materialize into responses. Rather than receiving a single, static reply, users might see a live preview of AI’s thinking, including key points, possible directions, or alternative solutions. They can then refine, select, or discard thoughts before finalizing the AI’s output.

### v2

@@ Abstract.
We envision the concept of Thoughtful AI, a new human–AI interaction paradigm in which the AI behaves as a continuously thinking entity. Unlike conventional AI systems that operate on a turn-based, input-output model, Thoughtful AI autonomously generates, iterates, and communicates its intermediate thought process throughout an interaction. In this position paper, we argue that this thoughtfulness unlocks new possibilities for human-AI interaction by enabling proactive AI behavior, facilitating continuous alignment with users, and fostering more dynamic interaction experiences. We outline the conceptual foundations of Thoughtful AI, illustrate its potential through example systems, and envision how this paradigm can transform human-AI interaction in the future.
@@ 1. Introduction
In 1950, Alan Turing famously asked, “Can machines think?” (Turing, 1950) This question has inspired decades of research in artificial intelligence (AI), with efforts ranging from symbolic reasoning and expert systems to today’s large language models (LLMs). Early philosophical and computational models of cognition viewed thought as a defining characteristic of intelligence—whether in humans or machines (Descartes, 1996; Newell, 1972; Miller and Cohen, 2001; Raichle et al., 2001). Recent LLMs have reignited discussions about what it means for an AI to “think”. Techniques such as Chain-of-Thought prompting (Wei et al., 2022) demonstrate that generating intermediate reasoning steps can significantly enhance model performance on complex tasks. Building on this foundation, models like OpenAI o1 (Jaech et al., 2024), Gemini Flash Thinking (Google DeepMind, 2025), and DeepSeek R1 (Guo et al., 2025) integrate reasoning into their training processes. These developments mark substantial progress in AI’s ability to problem-solve.
However, despite these advancements, current human-AI interaction paradigms remain fundamentally constrained. Existing systems still operate within a turn-based, input-output framework: Users issue a prompt, the AI generates a single response, and the cycle repeats. This limitation becomes evident in everyday AI interactions. Imagine a user asking ChatGPT: “Help me plan a surprise birthday party for my friend.” ChatGPT would generate a static response—perhaps listing general steps like choosing a venue, sending invitations, and selecting a menu. At this point, the AI has already stopped processing and just passively waits for the user’s next input. If the user later realizes they need a budget-friendly vegan restaurant for six people in New York, they must start a new turn, manually refining their query.
In contrast, consider instead how humans naturally collaborate during a brainstorming session. Each participant holds an internal train of thought that continuously updates in the background, even while someone else is speaking. In parallel, they may externalize immature, intermediate thoughts by sketching or annotating on a whiteboard, inviting immediate refinements and comments from others. This thinking process does not pause in discrete cycles: conversation flows as people update, erase, and reorganize ideas in real time.
@@
At the core of Thoughtful AI is to view AI’s thought as a first-class citizen in human-AI interaction: it is not merely hidden backend computation steps but a fundamental modality that contribute to AI’s capabilities to interact with humans. We identify four key traits that define this paradigm:
@@ 2. What makes AI “Thoughtful”?
Discussions around “thought” in LLM research often focus on a chain-of-thought approach, where a model generates intermediate reasoning steps to improve performance on reasoning tasks. From an HCI perspective, however, thought can be understood more broadly. We define Thoughtfulness as:
@@
Thoughtfulness refers to a system’s ability to continuously generate, develop, and selectively communicate its intermediate processes and responses over the course of an interaction.
@@ 2.1.1. Intermediate Medium
In most existing AI systems, users see only a final answer. This one-shot approach is limiting: once the user receives an answer, any clarifications or follow-up requests often require multiple back-and-forth queries. If the AI’s thinking was flawed or incomplete, the user has no window into how or why the system arrived at its conclusion.
@@ 2.2. Implications for HCI
The four traits of Thoughtful AI—an intermediate medium, full-duplex process, intrinsic driver, and a shared cognitive space—open up new possibilities for how humans interact with AI. Below, we discuss four implications for Human-AI Interaction.
@@ 2.2.2. Continuous Cognitive Alignment
2.2.3. Beyond Turn-based Interaction
Thoughtful AI prompts us to reconsider the fundamental structure of human–AI interaction. Traditional chatbot interfaces operate in discrete, back-and-forth messages, with the AI effectively going idle after sending each response. By contrast, a full-duplex process (Trait 2) enables the AI to continuously think and listen, even when the user is not actively providing input. This opens the door to interaction models that transcend simple chat boxes. For instance, an AI planning assistant could silently update its suggestions as it overhears new constraints in a virtual meeting, intervening only when necessary (Traits 2 and 3). More interestingly, AI and users can build a shared cognitive space (Trait 4), adding thoughts, annotations, and partial ideas to a collective interface. Rather than a linear log of exchanges, this collaborative canvas allows ideas to branch, merge, and evolve continuously—resembling a dynamic mind-map or sketchnote more than a turn-based chat.
2.2.4. Messy, Fragmented and Informal Interaction
Traditional AI systems demand users structure their inputs as complete, well-formed queries or commands. This formality creates friction: human thinking is inherently non-linear, often involving half-formed ideas and abrupt shifts in focus. Thoughtful AI embraces this messiness by using intermediate thoughts as the first-class citizen of interaction. This shift towards a more natural, fragmented, and informal interaction style: AI thinking can be messy, incomplete, and in-progress. In addition, instead of requiring users to provide polished, fully-formed queries or responses, users can also input partial ideas, rough thoughts, or even keywords, and the system can process and respond to these in a similarly fragmented manner. This approach mirrors how humans often think and communicate: by iterating on ideas, refining them over time, and bouncing off incomplete or tentative thoughts.
@@ 3. Example Projects
To further illustrate the practical applications of Thoughtful AI, we present two projects that embody its core principles. The first, Inner Thoughts, explores how AI can proactively engage in conversations by developing and evaluating its internal thoughts before participating. The second, ThinkaloudLM, investigates AI-generated thoughts as an interface component that allows users to observe and interact with in real-time.
@@ 3.1. Inner Thoughts
3.2. ThinkaloudLM
Figure 3. ThinkaloudLM: AI generates intermediate, fragmented thoughts in parellel to user input.
While Inner Thoughts explores AI’s ability to think internally, ThinkaloudLM investigates a complementary question: what if users could actively observe and engage with AI’s thought process?
@@ 3.2. ThinkaloudLM
By contrast, ThinkaloudLM envisions a system where users can engage with AI-generated ideas before they fully materialize into responses. Rather than receiving a single, static reply, users might see a live preview of AI’s thinking, including key points, possible directions, or alternative solutions. They can then refine, select, or discard thoughts before finalizing the AI’s output.